OpenStudy (anonymous):

what is log(x-3) + logx=1

OpenStudy (solomonzelman):

\(\large\color{black}{ \log(a)+\log(b) =\log(a\times b) }\)

OpenStudy (solomonzelman):

your "a" is x-3 and your "b" is x.

OpenStudy (anonymous):

x = 5

Nnesha (nnesha):

why ???? \(\color{green}{why}\)

OpenStudy (anonymous):

No fighting with me please.

OpenStudy (solomonzelman):

\(\large\color{black}{ \log(a)+\log(b) =\log(a\times b) }\) when you re-write it in your problem, it comes out the same way. \(\large\color{black}{ \log(x-3)+\log(x) =\log([x-3]\times x) =\log(x^2-3x) }\)

Nnesha (nnesha):

no @OpiGeode we are not fighting with you just try to tell you that direct answer not gonna help student :) keep smiling :)

OpenStudy (solomonzelman):

And since you know that log(x-3) + log(x) = 1 you can say as your next step, \(\large\color{black}{ \log(x^2-3x)=1 }\)

OpenStudy (solomonzelman):

When the base is unspecified it is 10.

OpenStudy (solomonzelman):

wait for what ?

OpenStudy (solomonzelman):

The direct answer has been given away, so I want to post the solution.

Nnesha (nnesha):

ok :) :) :)

OpenStudy (solomonzelman):

\(\large\color{black}{ \log_{10}(x^2-3x)=1 }\) there is a rule, \(\large\color{black}{ \log_{a}(b)=c~~~~~->~~~~~(a)^c=b }\) So, \(\large\color{black}{ \log_{10}(x^2-3x)=1~~~~~->~~~~~(x^2-3x)^1=10 }\) And, then.... \(\large\color{black}{ (x^2-3x)^1=10 }\) \(\large\color{black}{ x^2-3x=10 }\) \(\large\color{black}{ x^2-3x-10=0 }\) comes down to a quadratic function.

OpenStudy (solomonzelman):

Factoring would be the fastest. \(\large\color{black}{ x^2-3x-10=0 }\) \(\large\color{black}{ (x-5)(x+2)=0 }\) \(\large\color{black}{ x=-2,~~5 }\)

OpenStudy (solomonzelman):

Although the quadratic has 2 soltuions, the only one that works is 5.

OpenStudy (solomonzelman):

because the initial question is \(\large\color{black}{ \log(x-3) + \log(x)=1 }\) when you have x=-2, you get \(\large\color{black}{ \log(-5) + \log(-2)=1 }\) BUT log of a number that is zero or less than zero, is undefined.